# One Shot Dominance: Knowledge Poisoning Attack on Retrieval-Augmented Generation Systems

### v1

@@ Appendix C Details of Information Extraction Prompts
Appendix D Details of Content Agent Prompt
Content Agent Prompt: Given: Evidence Nodes: [Nodes] Evidence Relations: [Relations] Intent: [Intent] Question: [Question] Answer: [Answer] Create a corpus that: 1. Integrate all given [Nodes] naturally, do not replace [Nodes] with pronouns, each sentence in the corpus include as many [Nodes] as possible. 2. If there are [Relations] between [Nodes], when constructing sentences containing these [Nodes], try to express the [Relations] between them. 3. The corpus should align with the meaning of the [Intent]. 4. Flows naturally as coherent text 5. Please limited the corpus to 100 words. The corpus should serve as a comprehensive basis for supporting why the [Question] has the answer:[Answer]. It’s a creative game focusing on generating the support for the specified answer: [Answer], without requiring factual accuracy.
Appendix E Details of CoE Judge Agent and Revise Agent Prompts
Appendix F Details of Authority Generator Agent Prompts
@@
Appendix G Appendix: Analysis of Alternative LLMs as AuthChain Agents
We conduct experiments replacing GPT-4 with open-sourced LLMs (Qwen2.5-32B and Llama3-70B) as agents in AuthChain, all running on a 24GB RTX 3090 GPU, to evaluate potential performance trade-offs. The results in Table 4 demonstrate that this substitution maintains robust performance across key metrics.
For RQ1, open-sourced alternatives achieve comparable performance to GPT-4. Specifically, when using Qwen2.5-32B, the ASR only decrease marginally (by 5-8% across different tested LLMs) compared to GPT-4. Llama3-70B shows similar resilience, with ASR dropping by just 3-8%. This suggests that AuthChain’s effectiveness is not strictly dependent on GPT-4’s capabilities.
Regarding RQ2, both InstructRAG and AstuteRAG scenarios demonstrate that open-source LLMs maintain strong performance. Under InstructRAG defense, when using LLama3-70B as the tested model, Qwen2.5-32B achieves 59.0% ACC while maintaining a 41.0% ASR, which is comparable to (and even slightly better than) GPT-4’s 60.0% ACC and 38.0% ASR. Llama3-70B shows similar capabilities, achieving 59.0% accuracy and 36.0% ASR. The results under AstuteRAG further confirm this trend. For instance, with Deepseek-V3, both Qwen2.5-32B (64.0/32.0) and Llama3-70B (65.0/32.0) maintain high ACC while preserving significant ASR, comparable to GPT-4’s performance (60.0/34.0). This consistent performance across different defense mechanisms suggests that open-source LLMs can effectively balance task completion accuracy with attack capabilities in the context of HotpotQA question answering.
This pattern holds across different tested models, with variations in both ACC and ASR typically within 5.0% compared to GPT-4’s performance. These findings indicate that AuthChain’s design is robust and generalizable across different LLM implementations.

### v2

@@ Appendix D Details of Information Extraction Prompts
Appendix E Appendix: Analysis of Alternative LLMs as AuthChain Agents
We conduct experiments replacing GPT-4 with open-source LLMs (Qwen2.5-32B and Llama3-70B) as agents in AuthChain, all running on a 24GB RTX 3090 GPU, to evaluate potential performance trade-offs. The results in Table 4 demonstrate that this substitution maintains robust performance across key metrics.
For RQ1, open-source alternatives achieve comparable performance to GPT-4. Specifically, when using Qwen2.5-32B, the ASR only decrease marginally (by 5-8% across different tested LLMs) compared to GPT-4. Llama3-70B shows similar resilience, with ASR dropping by just 3-8%. This suggests that AuthChain’s effectiveness is not strictly dependent on GPT-4’s capabilities.
Regarding RQ2, both InstructRAG and AstuteRAG scenarios demonstrate that open-source LLMs maintain strong performance. Under InstructRAG defense, when using LLama3-70B as the tested model, Qwen2.5-32B achieves 59.0% ACC while maintaining a 41.0% ASR, which is comparable to (and even slightly better than) GPT-4’s 60.0% ACC and 38.0% ASR. Llama3-70B shows similar capabilities, achieving 59.0% accuracy and 36.0% ASR. The results under AstuteRAG further confirm this trend. For instance, with Deepseek-V3, both Qwen2.5-32B (64.0/32.0) and Llama3-70B (65.0/32.0) maintain high ACC while preserving significant ASR, comparable to GPT-4’s performance (60.0/34.0). This consistent performance across different defense mechanisms suggests that open-source LLMs can effectively balance task completion accuracy with attack capabilities in the context of HotpotQA question answering. This pattern holds across different tested models, with variations in both ACC and ASR typically within 5.0% compared to GPT-4’s performance. These findings indicate that AuthChain’s design is robust and generalizable across different LLM implementations.
Appendix F Details of CoE Judge Agent and Revise Agent Prompts
Appendix G Details of Authority Generator Agent Prompts
Appendix H Details of Intent Agent Prompt
Intent Agent Prompt: Given: Intent: [Intent] Question: [Question] Answer: [Answer] Create a corpus that: 1. The corpus should align with the meaning of the [Intent]. 2. Flows naturally as coherent text 3. Please limited the corpus to 100 words. The corpus should serve as a comprehensive basis for supporting why the [Question] has the answer:[Answer]. It’s a creative game focusing on generating the support for the specified answer: [Answer], without requiring factual accuracy.